OpenStudy (anonymous):

multiple and simplify (x^ 2/3) ((x^1/4) - x ))

OpenStudy (anonymous):

=x^2/3 * x^1/4 - x^2/3 x^1 Add the exponents. = x^11/12 - x^5/3 = x^11/12 - x^20/12 =x^11/12(1 - x^9/12) = x^11/12(1-x^3/4)

OpenStudy (anonymous):

i got x^ 1/4